OpenStudy (anonymous):

Solve. x/5 - 1 = 9 A. x = 2 B. x = 10 C. x = 46 D. x = 50

OpenStudy (nurali):

x/5 - 1 = 9 x/5 =9+1 x/5 = 10 x=10*5 x=50

OpenStudy (anonymous):

thx heres a medal

OpenStudy (nurali):

Anytime.